OpenStudy (anonymous):

HELP!!!! ill give medals and fan! What percentage of total students chose skateboarding? 10 20 30 none What percentage of boys chose walking? 30 25 45 none There are 20 students total 3 kids chose skateboarding and 3 boys chose walking I know 3 percent of 20 is 6 percent but that's not an answer and both answers have to be the same because there the same numbers. Its either none or 30% because thirty percent is the only answer out of the questions that are the same other than none.

OpenStudy (anonymous):

okay I found the answer for question two because it says what percentage of boys not total students and there is 10 boys and 3 chose walking so 30% I still need an answer for #1

OpenStudy (anonymous):

The percentage of skateboarders: \[ \frac{3}{20}*100\] keep in mind there is an option none

OpenStudy (anonymous):

the answer is 15 right? so it would be none?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

can you help with something else? also percentages

OpenStudy (anonymous):

ok go for it

OpenStudy (anonymous):

ok there are 500 students and I need to find out how many kids are in football than basketball 10% are soccer 27% are basketball 25% are baseball 38% are football

OpenStudy (anonymous):

ok first lets convert percentages into decimals : Divide the percentage by 100 to get your answer .10 are soccer .27 are basketball .25 are baseball .38 are football Now we multiply the decimal by the total amount of kids to see how many kids have actually played each sport .10*500 =50 kids Soccor .27*500 = 135 kids in basketabll .25*500 = 125 kids in Baseball .38*500 = 190 kids in Football

OpenStudy (anonymous):

Now you gotta do the rest :)

OpenStudy (anonymous):

WHat???O.o I need help:/ I don't know how.

OpenStudy (anonymous):

Well now that you have all the information how many more boys are in football than basketball. Just subtract

OpenStudy (anonymous):

I guessed 190 The answer choices are 325 125 190 55

OpenStudy (anonymous):

ok so I was right. Im sorry the message where you explained just loaded Thanks for helping could I ask for help with one more thing?

OpenStudy (anonymous):

yup go for it

OpenStudy (anonymous):

Thank you again okay so there are students surveyed about a program 190 like it 135 thinks it unnecessary 220 want to run for it for next year. if we made a circle graph what would be the central angle of the amount of students who like the program

OpenStudy (anonymous):

Aye maybe you shld ask that question to others cuz I dont seem to get it

OpenStudy (anonymous):

ohh waitt maybe I do one sec

OpenStudy (anonymous):

forget I am not familiar with this

OpenStudy (anonymous):

okay ill close this question and ask this question thanks for all your help!!!

OpenStudy (anonymous):

Np :)